SDO is GO 19 May 2011, 18:37 UTC
(200 words excerpt, click title or image to see full post)
Starting at 1600 UT today, May 19, 2011, the AIA 304 and 94 Å channels will be unavailable. The decontamination heaters will be turned on to drive contamination off of the CCD in that telescope. Other channels will be unaffected.